BILLS UPSTREAM CANOE TRIP TOOK 2.7 HOURS LONGER THAN THE RETURN TRIP. IF HE AVERAGED 2 MPH ON THE UPSTREAM TRIP AND 5MPH ON THE RETURN TRIP, THEN HOW LONG WAS THE UPSTREAM TRIP?
Answer by tedian.caseres(1)

vf^2 = vi^2 + 2ad
a= (vf-vi)/t
a= (5 mph- 2 mph) / 2.7 h
a= 1.11 m/h^2
rearranging d =( vf^2-vf^2)/2a
d= (25 m2/h2 - 4 m2/h2) / 2 ( 1.11 m/h2)
d= (21 m2/h2)/ 2.22 m/h2
d= 9.45 m